OpenStudy (annie12m):

Frankie’s electric bill total is $48 and contains three charges: a charge for electricity used, 5% tax on the charge for electricity used, and $6.30 in fees. Which statements are true? Choose exactly two answers that are correct. A. The equation that represents the price, e, of the electricity is e + 0.05e + 6.30 = 48. B. The equation that represents the price, e, of the electricity is 0.05e + 6.30 = 48. C. The cost of the electricity before taxes or fees is $39.71. D. The cost of the electricity before taxes or fees is $834.

OpenStudy (annie12m):

@jabez177 @retirEEd @RhondaSommer @truwhovian11

RhondaSommer (rhondasommer):

what do you think?

OpenStudy (annie12m):

b and c

OpenStudy (janu16):

A. The equation that represents the price, e, of the electricity is e + 0.05e + 6.30 = 48 C. The cost of the electricity before taxes or fees is $39.71. e + 5% of e + 6.30 = 48 e + 0.05e + 6.30 = 48 1.05e = 48 - 6.30 1.05e = 41.70 e = 39.71 39.71 + 5%(39.71) + 6.30 = 48 39.71 + 1.99 + 6.30 = 48 48 = 48

OpenStudy (janu16):

@annie12m

OpenStudy (annie12m):

ty